Grade 11Physical Chemistry

Some spherical ballon each of volume 2l are to be filled with hydrogen gas at one atm and 27 degree Celsius from a cylinder of volume 4l . The pressure of h2 gas inside the cylinder is 20 atm at 127 degree Celsius . Find the number of balloons which can be filled with h2 at 1 atm pressure and 300k .

Devesh
6 Years agoGrade 11

1 Answer

Vikas TU
ApprovedApproved Tutor Answer6 Years ago
Dear student
V baloon = 2 l = 2000 cm^3  
V cyllinder = 4l = 4000 cm^3
P = 20 atm 
T = 400k 
P1V1/T1 = P2V2/T2 
20*4000/400 = 1*V2 /300
V2 = 60000 cm^3
Vh2 = 60000-4000 = 56000 cm^3
No of baloon filled = 56000/2000 = 28 
Hope this helps 
Good luck